OpenStudy (anonymous):

solve, 4x8 + 12x4 +36 =0

ganeshie8 (ganeshie8):

start by factoring 4

ganeshie8 (ganeshie8):

and notice that it doesnt have any sing changes, in f(x) or f(-x) so you would get all imaginary roots. be prepared for that

ganeshie8 (ganeshie8):

4x8 + 12x4 +36 =0 4(x8 + 3x4 +9) = 0 x8 + 3x4 + 9 = 0

ganeshie8 (ganeshie8):

plugin x4 = t, then u wud see a quadratic, apply quadratic formula and solve t

OpenStudy (anonymous):

thanks a lot ganeshie8,if you don't mind could you solve it i compare my answer with yours,pliz

ganeshie8 (ganeshie8):

rest is pretty straightforward, you can use wolfram to compare your answers

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

you're wlcme :) there are no real solutions, oly imaginary solutions you would get when u apply quadratic formula. good luck !